OpenStudy (anonymous):

the substitution method to solve the system of equations. x + 2y = 15 2x – 3y = 2

OpenStudy (oaktree):

Well, substitution just means get one of the variables alone, so that's easy from the first equation.

OpenStudy (oaktree):

x=15-2y

OpenStudy (oaktree):

Plug that into the first equation and 2(15-2y)-3y=2, 30-4y-3y=2, 7y=28, y=4.

OpenStudy (oaktree):

Plug that into any equation (I'll choose x=15-2y) and x=15-2(4)=15-8=7.

OpenStudy (oaktree):

So the solution is x=7, y=4.

OpenStudy (anonymous):

thank you so so very much!<3